# Implicit Subgoal Planning with Variational Autoencoders for Long-Horizon Sparse Reward Robotic Tasks

### v1

@@ V Methodology
Figure 3: The Variational Autoencoder-based Subgoal Inference (VAESI) algorithm consists of three main components: a VAE-based Subgoal Generator, a Hindsight Sampler, and a Value Selector. The subgoal generator takes as input the current state and the desired goal and outputs subgoals to accomplish the long-horizon task. The Value Selector and the Hindsight Sampler are utilized to ensure that the subgoals are optimal and feasible, respectively.
To generate a sequence of feasible and optimal subgoals, we introduce the Variational Autoencoder-based Subgoal Inference (VAESI) algorithm. As illustrated in Fig. 3, our algorithm comprises three components: the VAE-based Subgoal Generator, the Hindsight Sampler, and the Value Selector. Initially, we employ the VAE-based Subgoal Generator to generate subgoal candidates grounded on the current state and the final goal. In order to demonstrate the feasibility of the subgoals, we endeavor to minimize the deviation between the subgoals generated and those sampled using the Hindsight Sampler. Finally, we use the Value Selector to filter the subgoals with the highest state value to guarantee the subgoals are optimal.
V-A Variational Autoencoder-based Subgoal Generator
@@
Given an initial state $s_{0}$ and a final goal $g$, we expect that the robot can infer a series of subgoals $\{\hat{g}_{1},\hat{g}_{2},\cdots,\hat{g}_{n-1}\}$ to guide the robot to the final goal $g$. These subgoals will subsequently be utilized to bootstrap the low-level action policy $\pi$, thereby generating the actions to accomplish the subgoal. This approach provides a more intensive reward to the low-level strategy compared to the direct use of the final goal.
@@
Generally, subgoal generators are always represented by explicit feed-forward models, which map subgoal space directly from state space and goal space. Current research [20] in behavior cloning has demonstrated that implicit models possess a superior capacity to learn long-horizon tasks more effectively than their explicit counterparts. This evidence incites us to employ the implicit model in subgoal generators, which is anticipated to enhance the ability to solve long-horizon tasks significantly.
@@
Based on these insights, we propose a hybrid subgoal generator, which is fundamentally a Conditional Variational Autoencoder (CVAE) [38]. Within this framework, the encoder $\psi$ serves as an explicit model that can deduce a sequence of subgoals, whereas the decoder $\psi^{\prime}$ operates as an implicit model, specifically engineered to augment the accomplishment of the subgoals, as illustrated in Fig. 4. After preprocessing state $s$ and the final goal $g$, the encoder computes the mean $\mu$ and the standard deviation $\sigma$ via two fully connected networks. The subgoal is then sampled from the normal distribution modeled using these calculated values of $\mu$ and $\sigma$. The decoder takes the subgoal $\hat{g}_{i}$ as input and is conditioned on the current state $s_{0}^{i}$ to generate the reconstructed final goal $g^{\prime}$. Note that we also utilize the reparameterization trick to make the encoder and decoder differentiable.
@@
Unlike the original CVAE, where the latent space serves as a low-dimension mapping from inputs, we utilize the structure of the CVAE but interpret the latent space as the subgoal space. By modeling the conditional option policy $\psi$, which takes as inputs the initial state $s_{0}^{i}$ of option $i$ and the final goal $g$, we transform the subgoal generation approach from Initial Planner (which infers subgoals directly) to Incremental Planner, as shown in Fig. 1 Left. That is, the subgoals can be inferred incrementally by:
$\hat{g}_{1}$ $=\psi(s_{0}^{1},g)$ (6)
$\hat{g}_{2}$ $=\psi(s_{0}^{2},g)$
$\vdots$
$\hat{g}_{n-1}$ $=\psi(s_{0}^{n-1},g)$
where $s_{0}^{1}=s_{0}$. To guarantee the robustness of the inferred subgoals, and prevent the agent from remaining at one subgoal, we set a time limit of $T$. The option policy $\psi$ deduces a new subgoal either upon the robot reaching the current subgoal or upon the expiration of the time step allocated to the current subgoal, that is,
@@
where $s^{i-1}_{e}$ is the terminal state when the robot reaches the subgoal $\hat{g}_{i-1}$ of option $i-1$. We denote the implicit model as $\psi^{\prime}(g^{\prime}|s_{0}^{i},\hat{g}_{i})$, which accelerates the more accurate generation of subgoals by minimizing the distance between the estimated final goal $g^{\prime}$ and $g$. Thus, the objective of the VAE subgoal generator $\mathcal{L}_{VAE}$ can be formulated as
$\mathcal{L}_{VAE}=\mathbb{E}\left[\|g-\psi^{\prime}(g^{\prime}|s_{0}^{i},\hat{g}_{i})\|^{2}\right]$ (8)
Note that a Kullback-Leibler (KL) Divergence term $D_{KL}[\psi(\hat{g}|s_{0}^{i},g_{i})\|p(\hat{g}_{i})]$ is also present in the original VAE loss [39], where $p(\hat{g}_{i})$ is referred to as a prior and always set to $\mathcal{N}(0,I)$. An improved prior will be introduced in the following section.
@@ V-B Hindsight Sampler for Feasible Subgoals
Ensuring the feasibility of the generated subgoals is of importance for the following reasons. Firstly, certain subgoals adhere to strict execution orders, implying that a robot can only execute a specific subgoal once it has completed the necessary prerequisite tasks. Second, some subgoals might be unfeasible within the goal space. Hence, the subgoal inferred by the VAE Subgoal Generator must ensure that all subgoals exist in a valid state, i.e., each subgoal is within reach from its preceding subgoal, which ensures the feasibility of the entire subgoal sequence.
@@
We expect the distribution of generated subgoals will closely approximate the true valid subgoal distribution [27]. The KL-divergence between the distribution of the subgoal generator $\psi$, and the true valid subgoal distribution, denoted as $P_{valid}$ is expressed as follows,
$D_{KL}(\psi(\cdot|s,g)||P_{valid}(\cdot)).$ (9)
The primary challenge arises from the unknown subgoal distribution $P_{valid}$, which complicates the estimation of the KL divergence. To address this challenge, we present a Hindsight Sampler, an approach based in the Hindsight Experience Replay buffer [36], designed to sample valid subgoals from an offline dataset. The core idea is to regard the achieved goal reached at the end of the trajectory as the final goal and recalculate the rewards for all states in the current trajectory. This revised trajectory is then employed to train the option policy. As shown in Fig. 5, during the training period, we first sample the trajectory $\tau_{D}$ with total time step $T_{\tau}$ from the offline dataset $\mathcal{D}$, and take the final achieved goal of this trajectory as the desired goal $g$ of the current trajectory. We then select waypoints from the trajectory as subgoals at intervals of $T_{\tau}/n$, denoted as $\left\{\hat{g}_{1}^{\prime},\hat{g}_{2}^{\prime},\cdots,\hat{g}_{n-1}^{\prime}\right\}$.
@@
where $P_{\psi}(\hat{g}_{i}^{\prime})$ is the probability of sampled subgoals under the distribution of $\hat{g}_{i}$, which is generated by the VAE-based Subgoal Generator. Maximizing the log probability of $\hat{g}_{i}^{\prime}$ allows us to make the subgoals generated by the option strategy $\psi$ feasible. Note that the number of subgoals will only be determined during the training phase to divide the entire trajectory into smaller segments. There is no requirement to allocate it during the task execution.
Input: Option policy $\psi$, low-level policy $\pi$, state value function $V$, time limit $T$, distance function $D$, initial state $s_{0}$, final goal $g$
@@ V-C Value Selector for Optimal Subgoals
The Universal Value Function Approximator (UVFA) plays a critical role in goal-conditioned RL. It functions as a mapping from the goal-conditioned state space to a nonlinear value function approximator [22]. We denote the state value function as $V$. In the context of goal-conditioned RL, it is typically utilized to assess the current state under the condition of the goal. The value function $V$ of SAC in goal-conditioned RL involves the extra entropy bonuses from every time step utilized to assess the state, denoted as follows,
$\begin{split}V(s,g)=\mathbb{E}_{g\sim\rho_{g},\tau\sim\rho_{\tau}}&\bigg{[}\sum_{t}\gamma^{t}r(s_{t},a_{t},g)\bigg{.}\\ &+\alpha\mathcal{H}(\pi(\cdot|s_{t},g))\mid s_{0}=s\bigg{.}\bigg{]}\end{split}$ (11)
In our algorithm, $V$ serves a dual purpose: guides the updates of low-level policy and selects optimal subgoals for the high-level planner. To generate subgoal $\hat{g}$ at current state $s_{t}$, conditioned on the final goal $g$, we initially sample $K$ candidates $\left\{\hat{g}_{1},\hat{g}_{2},\cdots,\hat{g}_{k},\cdots,\hat{g}_{K}\right\}$ using the VAE-based Subgoal Generator, where $\hat{g}_{k}\sim\psi(s_{t},g)$. As shown in Fig. 3, the candidates are then ranked using the state value function $V$, where higher state values represent better subgoals. Hence, the subgoal state corresponding to the highest value will be selected as the optimal subgoal $\hat{g}^{*}$ for the action policy, that is,
$\hat{g}^{*}=\arg\max_{\hat{g}_{k}}\{V(s,\hat{g}_{k}),k\in[1,K]\}$ (12)
@@ 2 Sample subgoals $\hat{g}^{\prime}$ by using Hindsight Sampler
3 Calculate $\mathcal{L}_{VAE}\leftarrow$ (8), $\mathcal{L}_{HS}\leftarrow$ (10)
@@ 4 Update option policy by $\psi^{*},\psi^{\prime*}=\arg\min_{\psi,\psi^{\prime}}\mathcal{L}$
Algorithm 2 VAESI
Alg. 1 offers a detailed description of the process of collecting rollout experience for the experience replay buffer. The function $D(g_{1},g_{2})$ is employed to estimate the distance between goals $g_{1}$ and $g_{2}$. We will initially infer a set of potential subgoal candidates via the VAE-Subgoal Generator when the current subgoal is achieved or the time limit $T$ is expirated. Then, the optimal subgoal is selected using the Value Selector. By replacing the final goal with the selected subgoal continuously, we can obtain trajectories of transitions $(s_{t},a_{t},r_{t},s_{t+1},\hat{g}^{*})$ and store it to the replay buffer $\mathcal{D}$.
The entire training process of VAESI is shown in Alg. 2. At first, a mini-batch $\mathcal{B}$ is sampled from $\mathcal{D}$, consisting of $N-1$ transitions of $(s_{t},a_{t},r_{t},s_{t+1},g)$, where $t\in[1,N]$. For each transition, the final goal $g$ is relabeled by the future achieved goal, and valid subgoals $\hat{g}^{\prime}$ will be sampled by using the Hindsight Sampler. Then, the objective of the VAESI is formulated as:
$\mathcal{L}=\mathcal{L}_{VAE}+\beta\mathcal{L}_{HS}$ (13)
@@ VI Experiments and Analysis
This section begins with an overview of the experimental task and setup. Then, we provide both quantitative and qualitative analyses of our method. Lastly, we conduct experiments in the real world to present the adaptability of our algorithm.
@@ VI-A Tasks
Our algorithms are implemented in four tasks: a locomotion task, 1) AntMaze, as shown in Fig. 2, and three manipulation tasks, 2) PushOverObstacle, 3) Stack, and 4) Push4, as shown in Fig. 6. The left and right columns present the simulation tasks and real-world tasks. All simulation environments are constructed utilizing the base of Gymnasium Robotics [41]. The dimensional information of the state space $\mathcal{S}$, action space $\mathcal{A}$, goal space $\mathcal{G}$, and the episode length $L$ are provided in Table I.
Figure 6: Three manipulation tasks, Stack, PushOverObstacle, and Push4 are considered to evaluate our algorithm. We mark the initial state and the desired goal as transparent spheres and solid cubes, respectively.
AntMaze AntMaze is a locomotion task performed in a U-shaped maze containing an 8-degree-of-freedom ant that tries to reach the final goal (the green sphere) through repeated trials. The ant is positively rewarded only when it reaches the final goal. The observation space has 26 axes and consists of positional states of different parts of the joints. The action space is the torque applied at the hinge joints [42].
PushOverObstacle Ren et al. [34] first propose the PushOverObstacle task, which is an extension of the Push task, where the initial state (green squares) and desired goal (green sphere) are uniformly spawned on different sides of an obstacle. The obstacle is a static wall that cannot be moved and is marked in red. The robot’s goal is to bypass the obstacle to reach the desired position.
Stack The block stacking task consists of picking up three blocks and placing them in sequence at the target location. The dimension of the observation space is 55, containing information about the fixture blocks. The action space contains the position of the fixture end-effector and the grasping signal. In addition, the robot receives no positive reward in the sparse reward setting until the entire task is completed. Multi-step planning makes it more difficult because we have to complete the whole task without disturbing all the completed subtasks [10].
Push4 This task combines the four block-pushing tasks of D4RL [42]. The observation space is 70-dimensional and consists of information about fixtures and blocks. The action space is the position of its end-effector and the gripper status. The robot must sequentially push four objects to the four target positions on the table.
@@
Parameters AntMaze PushOverObstacle Stack Push4
$\mathcal{S}$ 26 25 55 70
$\mathcal{G}$ 26 3 9 12
$\mathcal{A}$ 8 4 4 4
$L$ 100 100 300 400

### v2

@@ V Methodology
Figure 3: The Explicit-Implicit Subgoal Planning (EISP) algorithm consists of three main components: a Hybrid Subgoal Generator, a Hindsight Sampler, and a Value Selector. The subgoal generator takes as input the current state and the desired goal and outputs subgoals to accomplish the long-horizon task. The Value Selector and the Hindsight Sampler are utilized to ensure that the subgoals are optimal and feasible, respectively.
As illustrated in Fig. 3, our EISP algorithm comprises three components: the Hybrid Subgoal Generator, the Hindsight Sampler, and the Value Selector.
V-A Hybrid Subgoal Generator
@@
Given an initial state $s_{0}$ and a final goal $g$, we expect the robot to infer a series of subgoals $\{\hat{g}_{1},\hat{g}_{2},\cdots,\hat{g}_{n-1}\}$ to guide it to the final goal $g$. These subgoals are subsequently utilized to bootstrap the low-level action policy $\pi$, thereby generating the actions to accomplish each subgoal.
@@
Generally, most current subgoal generators [2, 24, 15] employ an explicit feed-forward model that maps subgoals directly from the state and the final goal. These methods are inspired by the human approach to solving long-horizon tasks, wherein humans generate subgoals to decompose long-horizon tasks into smaller, more manageable ones based on the final goal and the current state. However, one critical aspect often overlooked is that humans can also predict long-term outcomes using short-term subgoals conditioned on the current state. Current research [18] in behavior cloning has demonstrated that implicit models possess a superior capacity to learn long-horizon tasks more effectively than their explicit counterparts. This evidence incites us to employ the implicit model in subgoal generators, which is anticipated to enhance the ability to solve long-horizon tasks significantly.
@@
Based on these insights, we propose a hybrid subgoal generator, fundamentally structured as a Conditional Variational Autoencoder (CVAE) [36]. Within this framework, the encoder $\psi$ serves as an explicit model that can infer a sequence of subgoals based on the current state and the final goal, while the decoder $\psi^{\prime}$ functions as an implicit model, designed to reconstruct the final goal using the current state and the given subgoals, as depicted in Fig. 4. After preprocessing the state $s$ and the final goal $g$ through several fully connected layers, the encoder computes the mean $\mu$ and the standard deviation $\sigma$ via two fully connected networks. The subgoal is then sampled from the normal distribution modeled using these calculated values of $\mu$ and $\sigma$. The decoder takes the subgoal $\hat{g}_{i}$ as input and conditioned on the current state $s_{0}^{i}$ to generate the reconstructed final goal $g^{\prime}$. Notably, we also employ the reparameterization trick to make the encoder and decoder differentiable.
@@
Unlike the original CVAE, where the latent space serves as a low-dimensional mapping from the inputs, we employ the CVAE structure but interpret the latent space as the subgoal space. This approach distinguishes our work from previous studies [2], which also use the CVAE structure to generate subgoals but focus on reconstructing the subgoal using a learned latent representation of the transitions. By modeling the conditional option policy $\psi$, which takes as inputs the initial state $s_{0}^{i}$ of option $i$ and the final goal $g$, we transition the subgoal generation approach from the initial planner (which infers subgoals only at the initial state) to an incremental Planner (which infers subgoals based on current state). That is, the subgoals can be inferred incrementally by:
$\hat{g}_{i}=\psi(s_{0}^{i},g),\quad\text{for }i=1,\ldots,n-1$ (4)
where $s_{0}^{1}=s_{0}$. To ensure the robustness of the inferred subgoals and prevent the agent from lingering at one subgoal, we impose a time limit of $T$. The option policy $\psi$ deduces a new subgoal either when the agent reaches the current subgoal or when the allocated time step for the current subgoal expires. This can be expressed as follows:
@@
where $s^{i-1}_{e}$ is the terminal state when the robot reaches the subgoal $\hat{g}_{i-1}$ of option $i-1$. We denote the implicit model as $\psi^{\prime}(g^{\prime}|s_{0}^{i},\hat{g}_{i})$, which is designed to accurately reconstruct the final goal from the subgoal. The objective is to maximize the log probability of the final goal, $\log{p_{\psi^{\prime}}(g)}$, by maximizing its evidence lower bound (ELBO) [37], formulated as:
$\mathbb{E}_{p_{\psi}(\hat{g}|g,s_{0}^{i})}\Big{[}\log\frac{p_{\psi^{\prime}}(g,\hat{g},s_{0}^{i})}{p_{\psi}(\hat{g}|g,s_{0}^{i})}\Big{]}$ (6)
Thus, the objective of the hybrid subgoal generator $\mathcal{L}_{HB}$ can be defined as follows:
$\mathbb{E}_{p_{\psi}(\hat{g}_{i}|g,s_{0}^{i})}\big{[}\log p_{\psi^{\prime}}(g|\hat{g}_{i},s_{0}^{i})\big{]}-D_{KL}\big{[}p_{\psi}(\hat{g}_{i}|s_{0}^{i},g)\|p(\hat{g}_{i})\big{]}$ (7)
where $p(\hat{g}_{i})$ is the prior distribution over the subgoals, typically a standard normal distribution $\mathcal{N}(0,I)$. $D_{KL}[\psi(\hat{g}|s_{0}^{i},g_{i})\|p(\hat{g}_{i})]$ denotes the Kullback-Leibler (KL) Divergence [37] between the prior distribution $p(\hat{g}_{i})$ and the distribution of the generated subgoals by $\psi$. A detailed derivation can be found in Appendix -B. We will introduce an improved prior distribution for the subgoals in the following section.
@@ V-B Hindsight Sampler for Feasible Subgoals
The subgoals sampled from the ground truth subgoal distribution ensure that the transition probability from the current subgoal $\hat{g}_{i-1}$ to the next subgoal $\hat{g}_{i}$ is greater than zero. Thus, the distribution of the generated subgoals should closely approximate the true subgoal distribution, which can be achieved by minimizing the KL divergence between the distribution of the generated subgoals by $\psi$ and the true valid subgoal distribution $p_{t}(\hat{g}_{i})$. However, the primary challenge arises from the unknown ground truth subgoal distribution $p_{t}(\hat{g}_{i})$, which complicates the estimation of the KL divergence.
Empirically, the subgoal distribution within the offline dataset $\mathcal{D}$ can be approximately regarded as the ground truth subgoal distribution. Therefore, we adopt the subgoal distribution of the offline dataset as our prior distribution and aim to minimize the KL divergence between them. To this end, we present a Hindsight Sampler, an approach based on the Hindsight Experience Replay buffer [34], designed to sample valid subgoals from an offline dataset. The core idea is to regard the achieved goal reached at the end of the trajectory as the final goal and recalculate the rewards for all states in the current trajectory. This revised trajectory is then employed to train the option policy. As shown in Fig. 5, during the training period, we first sample the trajectory $\tau_{D}$ with total time step $T_{\tau}$ from the offline dataset $\mathcal{D}$, and take the final achieved goal of this trajectory as the desired goal $g$ of the current trajectory. We then select waypoints from the trajectory as subgoals at intervals of $T_{\tau}/n$, denoted as $\left\{\hat{g}_{1}^{\prime},\hat{g}_{2}^{\prime},\cdots,\hat{g}_{n-1}^{\prime}\right\}$.
@@
where $p_{\psi}(\hat{g}_{i}^{\prime})$ is the probability of sampled subgoals under the distribution of subgoals generated by the hybrid subgoal generator.
Input: Option policy $\psi$, low-level policy $\pi$, state value function $V$, time limit $T$, distance function $Dist$, initial state $s_{0}$, final goal $g$
@@ V-C Value Selector for Optimal Subgoals
The Universal Value Function Approximator (UVFA) plays a critical role in goal-conditioned RL. It functions as a mapping from the goal-conditioned state space to a nonlinear value function approximator [39]. We denote the state value function as $V$. In the context of goal-conditioned RL, it is typically utilized to assess the current state under the condition of the goal. The value function $V$ of SAC in goal-conditioned RL involves the extra entropy bonuses from every time step utilized to assess the state, denoted as follows:
$V(s,g)=\mathbb{E}_{g\sim\rho_{g},\tau\sim\rho_{\tau}}\Big{[}\sum_{t}\gamma^{t}r(s_{t},a_{t},g)+\alpha\mathcal{H}(\pi(\cdot|s_{t},g))\Big{]}$ (9)
where $s_{0}=s$. In our algorithm, $V$ serves a dual purpose: guides the updates of low-level policy and selects optimal subgoals for the high-level planner. To generate subgoal $\hat{g}$ at current state $s_{t}$, conditioned on the final goal $g$, we initially sample $K$ candidates $\left\{\hat{g}_{1},\hat{g}^{2},\cdots,\hat{g}^{k},\cdots,\hat{g}^{K}\right\}$ using the Hybrid Subgoal Generator, where $\hat{g}^{k}\sim\psi(s_{t},g)$. As shown in Fig. 3, the candidates are then ranked using the state value function $V$, where higher state values represent better subgoals. Hence, the subgoal state corresponding to the highest value will be selected as the optimal subgoal $\hat{g}^{*}$ for the action policy, that is,
$\hat{g}^{*}=\arg\max_{\hat{g}^{k}}\{V(s,\hat{g}^{k}),k\in[1,K]\}$ (10)
@@ 2 Sample subgoals $\hat{g}^{\prime}$ by using Hindsight Sampler
3 Calculate $\mathcal{L}_{HY}\leftarrow$ (7), $\mathcal{L}_{HS}\leftarrow$ (8)
@@ 4 Update option policy by $\psi^{*},\psi^{\prime*}=\arg\min_{\psi,\psi^{\prime}}\mathcal{L}$
Algorithm 2 EISP
Alg. 1 offers a detailed description of the process of collecting rollout experience for the experience replay buffer. The function $D(g_{1},g_{2})$ is employed to estimate the distance between goals $g_{1}$ and $g_{2}$. We will initially infer a set of potential subgoal candidates via the Hybrid Subgoal Generator when the current subgoal is achieved or the time limit $T$ is expirated. Then, the optimal subgoal is selected using the Value Selector. By replacing the final goal with the selected subgoal continuously, we can obtain trajectories of transitions $(s_{t},a_{t},r_{t},s_{t+1},\hat{g}^{*})$ and store it to the replay buffer $\mathcal{D}$.
The entire training process of EISP is shown in Alg. 2. At first, a mini-batch $\mathcal{B}$ is sampled from $\mathcal{D}$, consisting of $N-1$ transitions of $(s_{t},a_{t},r_{t},s_{t+1},g)$, where $t\in[1,N]$. For each transition, the final goal $g$ is relabeled by the future achieved goal, and valid subgoals $\hat{g}^{\prime}$ will be sampled by using the Hindsight Sampler. Then, the objective of the EISP is formulated as:
$\mathcal{L}=\mathcal{L}_{HY}+\beta\mathcal{L}_{HS}$ (11)
@@ VI Experiments and Analysis
Figure 6: Four manipulation tasks, i.e., Stack, Push, OpenDrawer, and Store. For the Stack and Push tasks, we label the initial states and the desired goals as solid cubes and transparent spheres. For the OpenDrawer and Store tasks, we use yellow dashed rectangles to represent the initial position of the blocks.
@@ VI-A Tasks
We evaluate EISP in four tasks: 1) Stack, 2) Push, 3) OpenDrawer, and 4) Store, as depicted in Fig. 6. The left and right columns present the simulation tasks and real-world tasks. All simulation environments are created using the Gymnasium Robotics [40] Framework. The dimensional information of the state space $\mathcal{S}$, action space $\mathcal{A}$, goal space $\mathcal{G}$, and the episode length $L$ are provided in Table I.
Stack The block stacking task consists of picking up three blocks and placing them in sequence at the target location. The dimension of the observation space is 55, containing information about the fixture blocks. The action space contains the position of the fixture end-effector and the grasping signal. Multi-step planning makes it more difficult because we have to complete the whole task without disturbing all the completed subtasks [9].
Push4 This task combines the four block-pushing tasks of D4RL [41]. The observation space is 70-dimensional and consists of information about fixtures and blocks. The action space is the position of its end-effector and the gripper status. The robot must sequentially push four objects to the four target positions on the table. The situations can become more complex when some blocks are blocked by others, necessitating the robot to push the blocks in a specific order.
OpenDrawer The robot needs to open the drawer by first removing the block and then pulling the handle. We evaluate EISP in both the state-based and image-based environment. In the state-based environment, the observation space is 42-dimensional, containing information about the drawer, the handle, the block and the gripper. In the image-based environment, the observation space is a $3\times 48\times 48$ image. The action space is the position of the end-effector and the gripper status. The task is challenging due to it contains multiple skills (Pushing the block and pulling the handle) and task dependencies.
Store This task requires the robot to first open the drawer, then pick and place the block into the drawer, and finally close the drawer. Multiple skills like pushing, picking, placing and grasping are involved in this task. Like the OpenDrawer task, we also conduct experiments in both the state-based and image-based environment. The settings of the observation and action space are the same as the OpenDrawer task.
@@
Env Stack Push4
OpenDrawer
Store
$\mathcal{S}$ 55 70 42 42
$\mathcal{G}$ 9 12 6 6
$\mathcal{A}$ 4 4 4 4
$L$ 300 400 300 1000